OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360°.

OpenStudy (kohai):

First of all, do you know what a polar coordinate is?

OpenStudy (anonymous):

no im in an online class and i have to teach myself @kohai

OpenStudy (kohai):

So, do you know what quadrant this point is in?

OpenStudy (anonymous):

no

OpenStudy (kohai):



OpenStudy (anonymous):

ok

OpenStudy (kohai):

So do you know how to even plot this point...?

OpenStudy (kohai):

I'm really confused as to how you made it to trig when you can't do pre-algebra stuff...

OpenStudy (ipwnbunnies):

OO KILL EM.

OpenStudy (anonymous):

\[put~2=r~\cos \theta~and~-2=r ~\sin \theta \] square and add \[2^2+\left( -2 \right)^2=r^2\left( \cos ^2 \theta+\sin ^2 \theta \right),r^2=8,r=2\sqrt{2}\] divide \[\tan \theta =-1=-\tan \frac{ \pi }{ 4 }=\tan \frac{ - \pi }{ 4 }=\tan \left( \pi-\frac{ \pi }{ 4 } \right),\tan \left( 2 \pi-\frac{ \pi }{ 4 } \right)\] \[\theta=?\]